OpenStudy (anonymous):

I really need to know how to solve -46 + 4 = 14

OpenStudy (cj49):

-46+4=14 -46+4=-42 -42=14 14+42=56

OpenStudy (joannablackwelder):

This is a false statement. It is like saying 2=1.